OpenStudy (anonymous):

What is the distance between the points (-2,1) and (3,7)

OpenStudy (anonymous):

use Pythagorus,

OpenStudy (anonymous):

distance formula.

OpenStudy (anonymous):

is the answer 37

sam (.sam.):

no

sam (.sam.):

\[\sqrt{(x-x_1)^2+(y-y_1)^2}\]

OpenStudy (anonymous):

times running out

OpenStudy (anonymous):

89

OpenStudy (anonymous):

???

sam (.sam.):

answer is in square root form

sam (.sam.):

\[\sqrt{(x-x_1)^2+(y-y_1)^2}\] try substitute (-2,1) and (3,7) into the formula

OpenStudy (anonymous):

yes square root of what A.11 B.37 C.61 D.89

sam (.sam.):

\[\sqrt{(x-x_1)^2+(y-y_1)^2}\] \[\sqrt{(-2-3)^2+(1-7)^2}=?\]